how may ways can a club select a president, vice president, and secretary from a group of 5 people?
1
Expert's answer
2020-12-08T07:32:20-0500

If you were asked in how many ways you could pick president, vice president and secretary such that the 1st person: president, 2nd person: vice president, 3rd person: secretary, you would have:

"5P3 = \frac{5!}{(5-3)!} = \frac{5!}{2!} = 3 \times 4 \times 5 = 60 \;ways"

Answer: 60 ways